OpenStudy (anonymous):

Please Help Me!! Will Medal!! Juno deposited $750 in a savings account that earns 4% interest compounded annually. If she does not deposit or withdraw any more money, how much money will there be in the account after 13 years? Round your answer to the nearest dollar.

OpenStudy (campbell_st):

this is compound interest \[A = P \times (1 + \frac{r}{100})^t\] in your question P = 750, r = 4 and t = 13 substitute and the calculate the answer

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

about 1249?

OpenStudy (wolf1728):

The formula you need is: Total = Principal × ( 1 + Rate )^years $750 13 years 4% Total = $750.00 × ( 1 + .04 )^13 Total = $750.00 × ( 1.04 )^13 Total = $750.00 × 1.6650735073 Total = $1,248.81 Calculator to check your answer: http://www.1728.org/compint.htm

OpenStudy (campbell_st):

makes sense to me